# CCE: Sample Efficient Sparse Reward Policy Learning for Robotic Navigation via Confidence-Controlled Exploration

### v7

@@ I Introduction
Reinforcement learning (RL) [1] is an important component of learning in robotic systems in a wide variety of complex domains, including search-and-rescue [2], disaster relief [3], and planetary exploration [4] applications. However, traditional RL methods often hinge on the availability of a dense reward structure [5], which may rarely be available in complex, unpredictable environments [6, 7, 8, 9]. Thus, reward function sparsity [10] frequently arise when applying RL methods to robotics problems, such as robotic navigation, due to their simple to implement and does not require expert knowledge to design like dense reward structures. However, this simplicity comes with a sparse feedback signal to learn from, thus leading to a limited exploration of the environment, and ultimately, to a suboptimal policy [11]. Existing literature in RL for robotics in sparse rewards has utilized methods such as entropy regularization [12], reward shaping [13, 14, 15, 16], curiosity-driven exploration [5, 17]. However, these methods are regularization-based that change the reward optimization objective which can lead to instability in learning [6].
Another major challenge of existing works comes from the sample inefficiency of exploration in sparse rewards. This issue is significant in robotic systems, due to the simulation-intensive nature of RL methods and the computational expense of realistic robotics simulation platforms [18]. Thus, RL training typically requires finding an optimal policy in a constrained sample budget. In this work, we diverge from previous exploration paradigms to respect the practicality of limited training samples. We do so by controlling exploration, rather than encouraging it based on a novel connection between the uncertainty, or confidence, of a policy and policy gradient estimation.
@@
Main Contributions: We present Confidence-Controlled Exploration (CCE), a general technique for improving the sample efficiency of RL methods for robotics navigation tasks with minimal modifications to the existing state-of-the-art methods. We introduce using the policy confidence, measured by the average conditional policy entropy over all states to control exploration. We focus on balancing exploration with the policy entropy by adaptively changing a salient feature in most RL methods: trajectory length. In policy gradient algorithms, trajectory length is the number of interactions the robot has with the environment before each gradient update during training. Our novel perspective is that exploration can be adaptively controlled by the number of samples needed to estimate the gradient. For navigation, an interaction is any action the mobile robot takes and the resulting state in the environment and the reward it receives.
To design CCE, our novel approach connects the need for the policy to explore to the mixing time of the policy. The mixing time of a policy is the amount of time it takes for a robot to reach its limiting behavior under that policy. When mixing time is large, ensuring that trajectories are sufficiently long is important for obtaining accurate estimates of system behavior and algorithmic essentials, such as policy gradients [19]. When mixing time is short, on the other hand, ensuring trajectory lengths are proportional frees up effort for performing more rapid policy updates and environmental exploration. Unfortunately, mixing time is typically difficult to estimate in complex environments [20, 21], so proxy methods for determining trajectory length are needed. In this work, we utilize a previously unexplored connection between a policy’s mixing time and its policy entropy to formulate CCE. We increase the sample efficiency by adaptively tuning the trajectory length required to update the policy parameter, thus decreasing the number of samples needed to learn an optimal policy for navigation tasks. Our results include:
@@ II-B Mixing Time and Spectral Gap
Though mixing times are useful in determining trajectory lengths, they are difficult to estimate in practice. Therefore, an alternative is estimating a lower bound. As described in [49], one such lower bound of mixing time is the absolute spectral gap, or simply spectral gap, of the induced transition matrix $\mathbb{P}_{\pi_{\theta}}$, defined to be the difference between $1$ and the second largest eigenvalue, $\lambda^{2}_{\theta}$:
@@ III-A Key Insight: Policy Entropy as a Proxy for Mixing Time
Given a policy $\pi_{\theta}$, we would ideally like to base trajectory length for our PG updates on the spectral gap $1-\lambda_{\theta}^{2}$ that lower bounds $\tau_{mix}^{\theta}$. However, without prior knowledge of the transition kernel $\mathbb{P}$ of the underlying MDP, it is difficult to calculate the spectral gap of the induced transition matrix, $\mathbb{P}_{\pi_{\theta}}$. One possible workaround is to estimate the transition kernel, but this quickly becomes intractable due to the $\Omega(|\mathcal{S}||\mathcal{A}|)$ number of samples required for each estimate.
@@
Figure 2: [LEFT] Visualization of $25$-by-$25$ gridworld with $4$ walls (black) dividing it into equal regions. [RIGHT] Positive correlation between a policy’s entropy and the spectral gap of its induced transition matrix over two different gridworld environments. Policy entropy is a viable proxy for mixing time.
@@ III-B CCE: Confidence-Controlled Exploration
3:     calculate $H_{c}$ using 4 for discrete actions, 5 for continuous actions
4:     calculate $K_{c}$ using 6
@@ IV Experiments and Results
This section provides a detailed empirical study and comparison of the proposed CCE method in Algorithm 1.
Figure 3: Comparison of four different trajectory-length schemes on [LEFT] even and [RIGHT] uneven terrain navigation tasks in Unity-based outdoor robot simulator. For REINFORCE, PPO, and SAC, using CCE as the trajectory length scheme leads to a navigation policy with a shorter path length return than baselines with limited samples. The drawn trajectories are sample representations of the robot’s odometry data collected. The trajectories pictured are representative of all $3$ algorithms since we found that each trajectory-length scheme leads to similar navigation regardless of the RL algorithm. See Figure 4 for training details.
Figure 4: [TOP] Learning curves during training and [BOTTOM] percent of sample budget of $\sim 4.5e4$ samples during navigation tasks for the even terrain shown on the left of Figure 3. Comparison of [LEFT] REINFORCE, [MIDDLE] PPO, and [RIGHT] SAC, with constant and adaptive trajectory lengths. For each algorithm, CCE converges to a higher cumulative reward while expending less of the total budget. We ran $8$ to $10$ independent replications for each of the $12$ experiments. Although fixed trajectory length of $200$ saves the sample budget comparably to CCE, it achieves lower reward in all cases due to the robot flipping over thus causing the episode to end.
Figure 5: Sample navigation trajectories generated in simulated and real-world environments by policies trained with REINFORCE using CCE vs. fixed trajectory lengths at a fixed sample budget of $\sim 4.5e4$. We observed that the fixed trajectory length policies often fail at test time due to inefficient use of the available sample budget during training. In contrast, CCE policies successfully complete the tasks after training with the same limited sample budget. The experiments with a real Clearpath Husky robot validate that CCE can be transferred onto real robotic systems without significant performance degradation. The drawn trajectories are sample representations of the robot’s odometry data collected during each task.
RL Algorithms. We choose three RL algorithms to integrate CCE into for improved sample efficiency: REINFORCE, PPO, and SAC. As REINFORCE and PPO are both on-policy policy gradient methods and SAC is an off-policy method, we show that CCE can improve the sample efficiency of various classes of policy gradient algorithms. For PPO and SAC, we adaptively change the size of the replay buffer. For information on hyperparameters, such as network architecture, $K_{i}$, and $\alpha$, please see Table II in the Appendix 111we re-provide the link to the supplemental version for convenience https://arxiv.org/abs/2306.06192 of the supplemental version.
@@
Figure 6: Similar to Figure 4 but for uneven terrain shown on the right of Figure 3. CCE still consistently reaches the highest cumulative reward while saving the largest percentage of training data. As in Figure 4, although fixed trajectory length of $200$ saves the sample budget comparably to CCE, it achieves lower reward in all cases due to the robot flipping over thus causing the episode to end.
@@
  • $\alpha_{goal}\in[0,\pi]$- Angle between the robot’s heading direction and the goal;
@@
Inspired by [26], we define three reward functions to formulate the sparse rewards to train our navigation policies: Goal distance reward $r_{dist}$, Goal heading reward $r_{head}$, and Elevation reward $r_{elev}$ to formulate the sparse rewards to train our navigation policies. Hence,
$r_{dist}=\frac{\beta}{2}\mathcal{N}(\frac{d_{goal}}{2},\sigma^{2})+\beta\mathcal{N}(0,\sigma^{2}),$ (7)
$r_{head}=\mathbb{I}_{\{|\alpha_{goal}|\leq\pi/3\}},$ (8)
where, $\mathcal{N}$ denotes Gaussian distribution with variance $\sigma$ and $\beta$ is the amplitude parameter. We set the variance $\sigma=0.2$ to ensure that the $r_{dist}$ only includes two narrow reward peaks near the goal and half-way from the goal. Similarly, $r_{elev}$ is defined as $r_{elev}=\mathbb{I}_{\{|\phi|\geq\pi/6\}}\cup\mathbb{I}_{\{|\theta|\geq\pi/6\}}$. Our reward definitions are significantly sparse compared to the traditional dense rewards used in the literature [41].
(a) Even terrain navigation: The robot is placed in an obstacle-free even terrain environment and random goals in the range of 10-15 meters away from the starting position given at each episode. The state space is $s=[d_{goal},\alpha_{goal}]$ and the overall reward formulation $r_{even}=r_{dist}+r_{head}$.
(b) Uneven terrain navigation: The robot is placed in an obstacle-free, highly uneven terrain environment. To account for this complexity, we input the roll and pitch of the angle of the robot into the state space, $s=[d_{goal},\alpha_{goal},\phi,\theta]$. The objective is to navigate the robot to a goal location along relatively even terrain regions to avoid possible robot flip-overs. The reward is $r_{uneven}=r_{dist}+r_{head}-r_{elev}$.
@@
Table I: Navigation performance of the policies trained for even and uneven terrain navigation tasks with limited samples. SR denotes success rate, PL denotes average path length, and EC denotes elevation cost. Each method is tested for 100 trials in the outdoor simulator and the average values are reported in the table. Test environments are different from training environments.
IV-A Evaluation Metrics
@@ IV-B Results and Discussions
Policy Learning: We observe in Figure 4 and 6 that, for the outdoor robot simulator, CCE achieves similar or better reward returns with significantly fewer samples under both the navigation settings compared to the fixed and random trajectory lengths. Significantly, despite providing comparable returns to CCE, the sample efficiency of the constant and large trajectory lengths is significantly lower (see the bar plots in Fig. 4 and 6). For the agents using CCE, in the even terrain simulations, all agents started with an initial trajectory length of $180$, and in the uneven terrain, $200$. All those agents end with a final length of $300$.
Navigation Performance: We compare our method’s navigation performance qualitatively in Fig. 5 and quantitatively in Table I. To highlight the importance of sample efficiency during training, we use policies trained up to episode $150$ for a limited sample budget of $\sim 4.5e4$ to perform navigation tasks in even and uneven simulated terrains. The policies, except for ones from CCE, cannot consistently complete the navigation tasks, resulting in significantly lower success rates. Even successful trials of these policies lead to significantly longer paths with several loops instead of heading toward the goal. In contrast, CCE results in relatively straight and consistent paths toward the goals. We further observe that uneven terrain navigation is a comparatively challenging task even for traditional methods such as Ego-graph[51]. Hence, the success rates are relatively low due to the robot flip-overs that occurred during trials. However, CCE leads to higher success rates even compared to baseline methods such as ego-graph. We further observe that CCE results in lower elevation cost indicating that the generated paths are smoother in uneven terrain conditions. Finally, we deploy the policies trained with the outdoor simulator in the real world using a Clearpath Husky robot. With CCE, even and uneven terrain navigation policies can be transferred to the real world without significant performance degradation in terms of success rate, path length, and elevation cost while other policies show inconsistent performance similar to the simulations.
@@ V Conclusions, Limitations, and Future Works
An important limitation of the present work is the restricted scope of our experimental investigation of the connection between policy entropy and the induced spectral gap. Although we cannot explore the connection in complex environments due to spectral gap calculation impracticality, it can still merit investigation for simple environments with continuous state and action spaces. During real-world experiments, all the navigation policies often lead to high motor vibrations from infeasible velocities provided as the action. Hence, further investigation is required to generate dynamically feasible and smooth actions for better sim-to-real transfer.

### v8

@@ I Introduction
Reinforcement learning (RL) [1] is an important component of learning in robotic systems in a wide variety of complex domains, including search-and-rescue [2], disaster relief [3], and planetary exploration [4] applications. However, traditional RL methods often hinge on the availability of a dense reward structure [5], which may rarely be available in complex, unpredictable environments [6, 7, 8, 9]. Thus, reward function sparsity [10] frequently arise when applying RL methods to robotics problems, such as robotic navigation, due to its simplicity to implementation and lack of need for expert knowledge to design required dense reward structures. However, this simplicity comes with a sparse feedback signal to learn from, thus leading to a limited exploration of the environment, and to a premature exploitation of a suboptimal policy [11]. Existing literature in RL for robotics in sparse rewards has utilized methods such as entropy regularization [12], reward shaping [13, 14, 15, 16], curiosity-driven exploration [5, 17]. However, these methods are regularization-based that change the reward optimization objective which can lead to instability in learning [6].
Another major challenge of existing works comes from the sample inefficiency of exploration in sparse rewards. This issue is significant in robotic systems, due to the simulation-intensive nature of RL methods and the high computational computational associated with realistic robotics simulation platforms [18]. Thus, RL training typically requires finding an optimal policy in a constrained sample budget. In this work, we propose an exploration paradigm to respect the practicality of limited training samples. We achieve this by balancing exploration, rather than encouraging it through a novel connection between the parameterized policy’s confidence and its gradient estimation.
@@
Main Contributions: We present Confidence-Controlled Exploration (CCE), a general technique for improving the sample efficiency of RL methods for robotics navigation tasks with minimal modifications to the existing state-of-the-art methods. Our method leverages policy confidence, measured by the average conditional policy entropy across all states, to regulate exploration. Specifically, we focus on balancing exploration with policy entropy by adaptively adjusting a key feature in most RL methods: trajectory length, defined as the number of samples the robot collects in the environment before each gradient update during training. For navigation tasks, a data sample refers to any action the mobile robot takes at a given state, the resulting state transition in the environment, and the corresponding reward.
To design CCE, our novel approach connects the need for the policy to explore to the mixing time of the policy. The mixing time of a policy is the amount of time it takes for a robot to reach its limiting behavior under that policy. When mixing time is large, ensuring that trajectories are sufficiently long is important for obtaining accurate estimates of system behavior and algorithmic essentials, such as policy gradients [19]. Conversely, when mixing time is short, keeping trajectory lengths proportional to it frees up resources for faster policy updates and more efficient exploration of the environment. Unfortunately, mixing time is typically difficult to estimate in complex environments [20, 21], so proxy methods for determining trajectory length are needed. In this work, we utilize a previously unexplored connection between a policy’s mixing time and its policy entropy to formulate CCE. We increase the sample efficiency by adaptively tuning the trajectory length required to update the policy parameter, thus decreasing the number of samples needed to learn an optimal policy for navigation tasks. Our results include:
@@ II-B Mixing Time and Spectral Gap
Though mixing times are useful in determining trajectory lengths, they are difficult to estimate in practice. Therefore, an alternative is estimating a lower bound. One such quantity that can lower bound of mixing time is the absolute spectral gap, or simply spectral gap, of the induced transition matrix $\mathbb{P}_{\pi_{\theta}}$, $1-\lambda^{\theta}_{2}$, where $\lambda^{\theta}_{2}$ is the second largest eigenvalue of $\mathbb{P}_{\pi_{\theta}}$, as stated in the theorem below [49]:
Theorem II.1
Let $\tau_{mix}^{\theta}$ be the mixing time induced by policy $\pi_{\theta}$. Then,
$\tau_{mix}^{\theta}(\epsilon)\geq\mathcal{O}\left(\frac{1}{1-|\lambda^{\theta}_{2}|}\right).$ (4)
@@ III-A Key Insight: Policy Entropy as a Proxy for Mixing Time
Given a policy $\pi_{\theta}$, we would ideally like to base trajectory length for our PG updates on the spectral gap $1-\lambda^{\theta}_{2}$ that lower bounds $\tau_{mix}^{\theta}$. However, without prior knowledge of the transition kernel $\mathbb{P}$ of the underlying MDP, it is difficult to calculate the spectral gap of the induced transition matrix, $\mathbb{P}_{\pi_{\theta}}$. One possible workaround is to estimate the transition kernel, but this quickly becomes intractable due to the $\Omega(|\mathcal{S}||\mathcal{A}|)$ number of samples required for each estimate.
@@
Figure 2: [LEFT] Visualization of $25\times 25$ gridworld with $4$ walls (black) dividing it into equal regions. [RIGHT] Positive correlation between a policy’s entropy and the spectral gap of its induced transition matrix over two different gridworld environments. Policy entropy is a viable proxy for mixing time.
@@ III-B CCE: Confidence-Controlled Exploration
3:     calculate $H_{c}$ using 5 for discrete actions, 6 for continuous actions
4:     calculate $K_{c}$ using 7
@@ IV Experiments and Results
Figure 3: Comparison of four different trajectory-length schemes on [LEFT] even and [RIGHT] uneven terrain navigation tasks in Unity-based outdoor robot simulator. For REINFORCE, PPO, and SAC, using CCE as the trajectory length scheme for $150$ episodes leads to a policy with a shorter path length return than baselines with limited samples. The drawn trajectories are sample representations of the robot’s odometry data collected. The trajectories pictured are representative of all $3$ algorithms since we found that each trajectory-length scheme leads to similar navigation regardless of the RL algorithm. See Figures 4 and 6 for training details.
Figure 4: [TOP] Learning curves during training and [BOTTOM] percent of sample budget ($\sim 4.5e4$) saved during navigation tasks for the even terrain (left of Figure 3). Comparison of [LEFT] REINFORCE, [MIDDLE] PPO, and [RIGHT] SAC, with constant and adaptive trajectory lengths. For each algorithm, CCE converges to a higher cumulative reward while expending less of the total budget. We ran $8$ to $10$ independent replications for each of the $12$ experiments. Although a fixed trajectory length of $200$ saves the sample budget comparably to CCE, it achieves lower reward in all cases due to the robot flipping over thus causing episodes to end.
Figure 5: Sample navigation trajectories generated in simulated (a, b) and real-world (c, d) environments by policies trained with REINFORCE using CCE vs. fixed trajectory lengths at a fixed sample budget of $\sim 4.5e4$. We observed that the fixed trajectory length policies often fail at test time due to inefficient use of the available sample budget during training. In contrast, CCE policies successfully complete the tasks after training with the same limited sample budget. The experiments with a real Clearpath Husky robot validate that CCE can be transferred onto real robotic systems without significant performance degradation. The drawn trajectories are sample representations of the robot’s odometry data collected during each task.
RL Algorithms. We choose three RL algorithms to integrate CCE into for improved sample efficiency: REINFORCE, PPO, and SAC. As REINFORCE and PPO are both on-policy policy gradient methods and SAC is an off-policy method, we show that CCE can improve the sample efficiency of various classes of policy gradient algorithms. For PPO and SAC, we adaptively change the size of the replay buffer. For information on hyperparameters, such as network architecture, $K_{i}$, and $\alpha$, please see the Appendix.
@@
Figure 6: [TOP] Learning curves during training and [BOTTOM] percent of sample budget ($\sim 4.5e4$) saved during navigation tasks for the uneven terrain (right of Figure 3). Comparison of [LEFT] REINFORCE, [MIDDLE] PPO, and [RIGHT] SAC, with constant and adaptive trajectory lengths. For each algorithm, CCE converges to a higher cumulative reward while expending less of the total budget. We ran $8$ to $10$ independent replications for each of the $12$ experiments. Although a fixed trajectory length of $200$ saves the sample budget comparably to CCE, it achieves lower reward in all cases due to the robot flipping over thus causing episodes to end.
@@
  • $\zeta_{goal}\in[0,\pi]$- Angle between the robot’s heading direction and the goal;
@@
Inspired by [26], we define three reward functions to formulate the sparse rewards to train our navigation policies: Goal distance reward $r_{dist}$, Goal heading reward $r_{head}$, and Elevation reward $r_{elev}$ to formulate the sparse rewards to train our navigation policies. Hence, $r_{dist}=\frac{\beta}{2}\mathcal{N}(\frac{d_{goal}}{2},\sigma^{2})+\beta\mathcal{N}(0,\sigma^{2})$ and $r_{head}=\mathbb{I}_{\{|\zeta_{goal}|\leq\pi/3\}}$, where $\mathcal{N}$ denotes Gaussian distribution with variance $\sigma$ and $\beta$ is the amplitude parameter. We set the variance $\sigma=0.2$ to ensure that the $r_{dist}$ only includes two narrow reward peaks near the goal and half-way from the goal. Similarly, $r_{elev}$ is defined as $r_{elev}=\mathbb{I}_{\{|\phi|\geq\pi/6\}}\cup\mathbb{I}_{\{|\theta|\geq\pi/6\}}$. Our reward definitions are significantly sparse compared to the traditional dense rewards used in the literature [41].
(a) Even terrain navigation: The robot is placed in an obstacle-free even terrain environment and random goals in the range of 10-15 meters away from the starting position given at each episode. The state space is $s=[d_{goal},\zeta_{goal}]$ and the overall reward formulation $r_{even}=r_{dist}+r_{head}$.
(b) Uneven terrain navigation: The robot is placed in an obstacle-free, highly uneven terrain environment. To account for this complexity, we input the roll and pitch of the angle of the robot into the state space, $s=[d_{goal},\zeta_{goal},\phi,\theta]$. The objective is to navigate the robot to a goal location along relatively even terrain regions to avoid possible robot flip-overs. The reward is $r_{uneven}=r_{dist}+r_{head}-r_{elev}$.
@@
Table I: Navigation performance of the policies trained for even and uneven terrain navigation tasks with limited samples. SR denotes success rate, PL denotes average path length, and EC denotes elevation cost. Each method is tested for 100 trials in the outdoor simulator and the average values are reported in the table. CCE has a higher SR, lower PL, and lower EC than baseline approaches in the even terrain, and higher SR and lower EC in uneven. Test environments are different from training environments.
IV-A Evaluation Metrics for Navigation Paths Generated with Fixed Training Sample Budget
@@ IV-B Results and Discussions
Policy Learning: We observe in Figure 4 and 6 that, for the outdoor robot simulator, CCE achieves similar or better reward returns with significantly fewer samples under both the navigation settings compared to the fixed trajectory lengths and entropy regularization. Significantly, despite providing comparable returns to CCE, the sample efficiency of the constant and large trajectory lengths is significantly lower (see the bar plots in Fig. 4 and 6). For the agents using CCE, in the even terrain simulations, all agents started with an initial trajectory length of $180$, and in the uneven terrain, $200$. All those agents end with a final length of $300$.
Navigation Performance: We compare our method’s navigation performance qualitatively in Fig. 5 and quantitatively in Table I. To highlight the importance of sample efficiency during training, we use policies trained up to episode $150$ for a limited sample budget of $\sim 4.5e4$ to perform navigation tasks in even and uneven simulated terrains. The policies, except for ones from CCE, cannot consistently complete the navigation tasks, resulting in significantly lower success rates. Even successful trials of these policies lead to significantly longer paths with several loops instead of heading toward the goal. In contrast, CCE results in relatively straight and consistent paths toward the goals. We further observe that uneven terrain navigation is a comparatively challenging task even for traditional methods such as Ego-graph[51]. Hence, the success rates are relatively low due to the robot flip-overs that occurred during trials. However, CCE leads to higher success rates even compared to baseline methods such as ego-graph. We further observe that CCE results in lower elevation cost indicating that the generated paths are smoother in uneven terrain conditions. Finally, we deploy the policies trained with the outdoor simulator in the real world using a Clearpath Husky robot. With CCE, even and uneven terrain navigation policies can be transferred to the real world without significant performance degradation in success rate, path length, and elevation cost while other policies show inconsistent performance similar to the simulations.
@@ V Conclusions, Limitations, and Future Works
A limitation is the scope of our experimental investigation of the connection between policy entropy and the induced spectral gap. Although we cannot explore the connection in complex environments due to spectral gap calculation impracticality, it can still merit investigation for simple environments with continuous state and action spaces. In real-world experiments, policies often lead to high motor vibrations from infeasible velocities provided as actions. Hence, further investigation is required to generate dynamically feasible and smooth actions for better sim-to-real transfer.